Assume that when adults with smartphones are randomly selected, 47% use them in meetings or classes. If 10 adults smartphone users are randomly selected, find the probability that at least 3 of them use their smartphones in meetings or classes.
Answer #1

Binomial probability distribution formula: P(X) = nCx px qn-x

P(a person uses phone in meetings or classes), p = 0.47

q = 1 - p = 0.53

Sample size, n = 10

P(at least 3 of them use their smartphones in meetings or classes) = P(X 3)

= 1 - P(X < 3)

= 1 - P(0) - P(1) - P(2)

= 1 - 0.5310 - 10x0.47x0.539 - 10C2x0.472x0.538

= 1 - 0.00175 - 0.01551 - 0.06189

= 0.92085